40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL, Page 50 of 52  Not logged in ELOG logo
ID Date Author Type Category Subject
  142   Thu Jun 3 14:47:58 2010 FrankLaserLaserAligning AOM

Quote:

Quote:

 I'm aligning the AOM and maximizing the diffracted beam's power by positioning the AOM and adjusting the beam size by moving the lens.

For single pass, the maximum efficiency I could get is only ~60%, so for double pass, the power will be down to 36%, but for now I'll settle with this number.

I could not find the manual for Crystal technology AOM 3080-194. The closest one is model 3080-197 which is attached below.

I'm not sure what is the difference between the two model, but 3080-197 has 70% diffraction efficiency.

Because of adjusting the lens, the RefCav's beam path also changes, now I have to realign RefCav again.

 

  Another step for AOM alignment is adjusting the mirror that reflects the transmitted beam back to the AOM again.

The distance between the mirror and the center of the AOM should be the same as ROC of the mirror.

After this I should be able to start locking ACav.

 did you measure the power of the vco? How much is it if you tune it to maximum?

Here a copy of a general datasheet for the 3080-194. maximum efficiency is ~80% @2W RF power. You should ask peter about the detailed datasheet which comes with each AOM and contains measured values for the one you are using. Measured values depend on the beam size and RF power. Typical values are 87% in reality.

AOM3080-194.JPG

 Oh, I see, the beam diameter is 1100 um, I use 150um. I'll try changing the beam size and see what happens. Thanks Frank. I'll measure the power of the VCO too.

  141   Thu Jun 3 08:42:44 2010 FrankLaserLaserAligning AOM

Quote:

 I'm aligning the AOM and maximizing the diffracted beam's power by positioning the AOM and adjusting the beam size by moving the lens.

For single pass, the maximum efficiency I could get is only ~60%, so for double pass, the power will be down to 36%, but for now I'll settle with this number.

I could not find the manual for Crystal technology AOM 3080-194. The closest one is model 3080-197 which is attached below.

I'm not sure what is the difference between the two model, but 3080-197 has 70% diffraction efficiency.

Because of adjusting the lens, the RefCav's beam path also changes, now I have to realign RefCav again.

 

  Another step for AOM alignment is adjusting the mirror that reflects the transmitted beam back to the AOM again.

The distance between the mirror and the center of the AOM should be the same as ROC of the mirror.

After this I should be able to start locking ACav.

 did you measure the power of the vco? How much is it if you tune it to maximum?

Here a copy of a general datasheet for the 3080-194. maximum efficiency is ~80% @2W RF power. You should ask peter about the detailed datasheet which comes with each AOM and contains measured values for the one you are using. Measured values depend on the beam size and RF power. Typical values are 87% in reality.

AOM3080-194.JPG

  140   Thu Jun 3 00:52:37 2010 taracLaserLaserAligning AOM

 I'm aligning the AOM and maximizing the diffracted beam's power by positioning the AOM and adjusting the beam size by moving the lens.

For single pass, the maximum efficiency I could get is only ~60%, so for double pass, the power will be down to 36%, but for now I'll settle with this number.

I could not find the manual for Crystal technology AOM 3080-194. The closest one is model 3080-197 which is attached below.

I'm not sure what is the difference between the two model, but 3080-197 has 70% diffraction efficiency.

Because of adjusting the lens, the RefCav's beam path also changes, now I have to realign RefCav again.

 

  Another step for AOM alignment is adjusting the mirror that reflects the transmitted beam back to the AOM again.

The distance between the mirror and the center of the AOM should be the same as ROC of the mirror.

After this I should be able to start locking ACav.

Attachment 1: 97-02848-01r0.pdf
97-02848-01r0.pdf
  139   Tue Jun 1 19:26:41 2010 taracLaserLaserA mirror gets into the insulation's border

I'm aligning the AOM. The R=0.5m mirror's position crosses the insulation border by 1.5" (see attached picture.) The black line on the table shows the border of the insulator. The mirror is on a translational stage.

I'm thinking of 2 choices to solve this,

 1)using a mirror to turn the beam to the side of the table. The mirror will be placed after the AOM, around the edge the border.

2) using 2 mirrors (after the beam is split to RefCav and ACav's paths) to shift the beam path to the side of the table.

The first choice will be better, since I won't have to recalculate the mode matching, but there might be unexpected problems.

The VCO is working fine, I can see +/- 1st order beams coming out.

Attachment 1: Photo_37.jpg
Photo_37.jpg
  138   Tue Jun 1 14:16:00 2010 taracLaserLaserPSL setup: next step

This is the schematic for PSL setup.

http://131.215.115.52:8080/PSL_Lab/117

At this point, Pre Mode Cleaner (PMC) and Reference Cavity (RefCav) are locked. The rest will be locking Analyzer Cavity (ACav) and setting up for beat noise measurement.

ACav's beam path will have double pass AOM [Crystaltech 3080 194]. We'll use +1st order beam. When hook up the VCO, make sure that the power is on only when the VCO and the AOM are connected, otherwise the VCO dies.

Next is aligning the AOM. A good alignment will maximize the power of the +1st order beam. The beam should get close to the AOM's transducer as much as possible to minimize time delay.

The beam at the AOM will be focused to 75 um.

The mirror that reflects the beam back to the AOM is a 0.3m concave mirror, which will be placed 0.3 m away from the AOM. The reflected beam should completely overlap on itself. This will neutralize

the pointing instability when the modulating frequency shifts.

 

Then we can align ACav, this time I'll try not to remove the PMC when I scan the beam frequency (at~3-10Hz.) If the PMC cannot catch up with the laser, increase the gain of the PMC, sideband power. 

 ACav should be locked before Monday June 8.

 

  137   Tue Jun 1 07:39:47 2010 FrankLaserLaserRefCav is locked

Quote:

I see TEM00 transmitted beam out of RefCav. I think I have to fine tune the FSS gain a bit more becasue the spot still oscillates a bit.

The left monitor shows the spot from  PMC. The right monitor shows the spot from RefCav. it looks distorted becasue of the filter.

 

The common gain is 16.1 dB

slow actuator is 9 V

Fast Gain 15dB

Optimization is yet to be done. There is still plenty of reflected power.

 

 nice

  136   Sun May 30 22:49:28 2010 taracLaserLaserRefCav is locked

I see TEM00 transmitted beam out of RefCav. I think I have to fine tune the FSS gain a bit more becasue the spot still oscillates a bit.

The left monitor shows the spot from  PMC. The right monitor shows the spot from RefCav. it looks distorted becasue of the filter.

 

The common gain is 16.1 dB

slow actuator is 9 V

Fast Gain 15dB

Optimization is yet to be done. There is still plenty of reflected power.

 

Attachment 1: Photo_36.jpg
Photo_36.jpg
  135   Sat May 29 12:01:17 2010 FrankLaserLaserProgress on PSL: Aligning RefCav

Quote:

 Since the frequency of the laser going into RefCav is determined by PMC ,I decided to temporarily remove the PMC for now, so I can scan the laser frequency while aligning RefCav.

It might not be a good idea since the PMC might alter the beam path a little bit, but I just want to align the cavity first.

The plan is after  RefCav is aligned, I'll bring back the PMC and fine tune the beam going to RefCav again.

I still got many higher order modes coming out of the cavity.

 

 the PMC is locked to the laser, so it follows the NPRO frequency when you scan the frequency

  134   Fri May 28 20:46:14 2010 taracLaserLaserProgress on PSL: Aligning RefCav

 Since the frequency of the laser going into RefCav is determined by PMC ,I decided to temporarily remove the PMC for now, so I can scan the laser frequency while aligning RefCav.

It might not be a good idea since the PMC might alter the beam path a little bit, but I just want to align the cavity first.

The plan is after  RefCav is aligned, I'll bring back the PMC and fine tune the beam going to RefCav again.

I still got many higher order modes coming out of the cavity.

 

  133   Thu May 27 18:20:42 2010 taracLaserLaserAligning RefCav

I switched the cable from the 10W controller to the original controller for 100 mW laser. It is working well now, the cables are tied properly.

For now, I don't need to use the FSS servo card to scan the laser frequency.

I'm using a function generator for fast channel (PZT), and a voltage calibrator for slow channel (thermal control.)

The alignment is in progress. With the aid of a CCD camera and a macroscopic lens, looking for the beam position on the mirror is getting easier.

Currently I see some light at the back of the cavity.

  132   Thu May 27 11:18:51 2010 taracLaserLaserSchematic for FSS servo and Lightwave 126 1064 100mW

This is the schematic of PSL fss servo.

I have to make sure that the modulation voltage will not exceed the controller's limit (0-100V.)

 

Attachment 1: D980536-D.pdf
D980536-D.pdf
Attachment 2: npro125126_ds_cl_ae_020606-1.pdf
npro125126_ds_cl_ae_020606-1.pdf npro125126_ds_cl_ae_020606-1.pdf npro125126_ds_cl_ae_020606-1.pdf npro125126_ds_cl_ae_020606-1.pdf npro125126_ds_cl_ae_020606-1.pdf
  131   Wed May 26 21:11:00 2010 taracLaserLaserprogress on PSL setup

 

I add the broad band EOM in the beam path. After adjusting the periscope, I can steer the beam into the RefCav and see the reflected light. It's not aligned yet.

The 35.5 MHz is set on the table with a lens to focus the beam on the PD.

I'm not sure if I have to use the laser controller for 126 model or I could use the 10W laser to scan the beam, I'll consult Peter tomorrow.

Right now we are using the 10W laser controller to power the 100mW laser. The connector had been unstable, but now it's working fine.

It will be better if I can use the 10W controller to dither the laser frequency because I won't have to switch the cable, and avoid the risk of having to deal with the cable again.

 

 

Forgot to log this yesterday:

The PMC servo in medm's command window is correct.

 

I need to make SMA cables (properly insulated kind) too.

From 21.5 MHz PD to servo, 25 feet,

from 35.5 MHz PD to servo,   25 feet,

from 35.5 MHz EOM to signal box, 5 feet,

from 35.5 MHz LO to signal box, 20 feet.

  130   Tue May 25 17:01:57 2010 taracLaserLaserPMC cavity is locked

      After adding another 1/2 plate to have P wave into the cavity, I can lock the PMC cavity. It's been 30 minutes so far.

There is only one transmitted beam now (there were two when I used S wave.) 

 

     When I work or knock on the table, sometime the beam switches to another mode (might be it's side band.)

It's very close to the main TEM00 mode. I need to adjust the DC offset a little bit to get back.

 The incoming power is 9 mW, and the Transmitted beam has ~ 6 mW.

Current setting:

RF Amp: 7V

Phase shift: 2.87 + 180

Gain: 27.91 dB

DC offset: -2.3 V

   

  

     I tried to pre align the ref cavity. It's harder than I thought, can't see the beam that well. I'll have to check the manual for the laser controller, so that I can

scan the RefCav when I align the beam into the cavity.

  129   Mon May 24 19:40:12 2010 taracLaserLaserRef Cav prealignment

 

Dear elog

  I don't know why I can't lock the PMC. After changing the 21.5 MHz card because of the loosen SMA connector, the 21.5 MHz EOM is working.

The error signal looks good. I adjusted the gain, flipped the phase by 180 degree, and still cannot lock the cavity. 

(The medm was frozen this morning, Peter helped reset it back to work this after noon.) It might be insufficient amount of power coupling into the cavity.

The minimum reflected beam I could get is only ~1/2 of the total DC power, I'll try to align the beam and move the lens a little bit more to see if I can optimize it better.

So I skip this PMC part for now, and pre align the path to mode cleaner. The mode matching is ok. Two lens, f1=114.5 and f2=286.3 mm are good. A CCD behind the cavity is set in place.

 

  128   Mon May 24 19:30:15 2010 ranaLaserPMCPMC is almost ready to be locked

Quote:

You should make sure not to blow up the PMC PD: i.e. the total power on the PD out of lock should be less than 100 mW. The beam size on the PD should also be ~0.5 mm diameter with the actual beam waist being more like 0.3 mm dia.

PD should also be tilted by ~30 deg from normal incidence and the reflection dumped. Its OK if the RF output saturates somewhat at the peak of the PDH signal, but the in-lock RF level should be below ~50 mVrms into 50 Ohms.

 The power is small, our laser is 100 mW. The power on the PD is not saturated. Right now it's receiving ~10mW. It's tilted a bit and the reflected beam is blocked appropriately.

  127   Mon May 24 09:56:19 2010 ranaLaserPMCPMC is almost ready to be locked

You should make sure not to blow up the PMC PD: i.e. the total power on the PD out of lock should be less than 100 mW. The beam size on the PD should also be ~0.5 mm diameter with the actual beam waist being more like 0.3 mm dia.

PD should also be tilted by ~30 deg from normal incidence and the reflection dumped. Its OK if the RF output saturates somewhat at the peak of the PDH signal, but the in-lock RF level should be below ~50 mVrms into 50 Ohms.

  126   Thu May 20 21:50:35 2010 taracLaserLaserPMC is almost ready to be locked

 We still need two SMA cables. One connects between 21.5Mhz EOM and PMC servo card, another one connects between PMC PD and the servo card.

The 21.5 and 35.5 local oscillator were in the wrong slots, we fixed them.

The 21.5 MHz  photo diode that detects the reflected beam from PMC saturates at 15 mW.

Now I'm trying to optimize the PMC setup so that we have maximum transmittedd power.

  125   Wed May 19 00:18:45 2010 Tara ChalermsongsakLaserLaserPMC alignment

After checking the PMC servo card,

ramp signal goes to ext DC 

HV in is applied properly,

HV out is fixed.

Now the PMC is scanning and working fine.

The alignment is done. Although optimization is still needed,  I can work on the rest of the setup.

 

*note on PMC servo card

Ratio between Voltage input (Vin), V monitor (Vmon), and High Voltage output (HVout),

HVout = 24 Vin

Vmon = 1/50 HVout

Vmon = 24/50 Vin

 see the attachment

Attachment 1: D980352-B.pdf
D980352-B.pdf
  124   Tue May 18 15:21:12 2010 Tara ChalermsongsakLaserLaserPMC alignment

Quote:

I'm trying to align the PMC.  The transducer is connected to the PMC servo card's HV out.

HV in is driven by a function generator with triangular function. The output signal looks weird. It is not a nice triangular form, it's more like a u shape waveform connecting to each others with a plateau on top.

I'll check if the transducer on the PMC and the HV out from the card are working correctly or not. Right now, there is not a glimpse of signal coming out of the PMC on CCD.

don't the words HV IN ring a bell? look into the schematic. there are only a couple of inputs at the front so it's not too hard to figure out why there was already a BNC cable connected to that input. Did you check the monitor signal? If you would you would have seen that your HV supply is missing now

 

  123   Mon May 17 21:32:35 2010 Tara ChalermsongsakLaserLaserPMC alignment

I'm trying to align the PMC.  The transducer is connected to the PMC servo card's HV out.

HV in is driven by a function generator with triangular function. The output signal looks weird. It is not a nice triangular form, it's more like a u shape waveform connecting to each others with a plateau on top.

 

I'll check if the transducer on the PMC and the HV out from the card are working correctly or not. Right now, there is not a glimpse of signal coming out of the PMC on CCD.

  122   Thu May 13 20:56:57 2010 FrankLaserLaser100mW NPRO problems

we still have serious problems with the 100mW laser head. I traced it down to the connection between the PCB and the hermetically sealed optical part. There is a really loose connection somewhere. The connector seems to be OK, the PCB and all solder points are OK too (visually). But if you slightly touch the PCB you can see the yellow LED flickering, if you touch it a bit more the laser goes off and on. This happens too if you touch the D-SUB cable on the back. I added some little stress to the PCB when putting the thing back together, now the situation seems to be better, but is far away from being gone. In order to get started we are using it now as long as we can and think about a solution in the meantime. One option would be to fix the busted NPRO Peter has. This would probably take about a week or so. We have to align the laser diode and focusing lens and solder (!)  the focusing lens in place. The problem is that you can't align in vertical direction, so you have to remove the laser diode, put some more or less indium foil below and start again from zero. But it's an option. Peter had it already back to 550mW (out of 700mW) or so (for a couple of minutes, simply holding all parts in place. So the difficult part is to keep it permanent in the right place...

  121   Thu May 13 13:41:32 2010 Tara ChalermsongsakLaserLasermode matching from laser to PMC

Details for Mode Matching

 

1)   Laser to PMC

 

The laser has

              Wx = 155 um, 3.45 cm in front of the opening.

            Wy = 201 um, 2.8 cm in front of the opening.

 

The average number for calculation is w = 180 um, 3cm  in front of the opening.

First, we focus the beam to the EOM, w can be  250 – 500 um.

We pick 350 um.

F= 200 mm,

W1= 180 um

W2=350 um

Distance from w1 to the lens, d1, = 9.36”.

Distance from w1 to w2, L, = 22.8 “.

 

 

Then we mode match this beam to PMC

F1= 50.2mm

F2= 63mm

W1=350 um

W2= 370um

Distance from w1 to first lens, d1,         =157mm =6.2”

Distance from first lens to 2nd lens, d2 =120mm = 4.7”

Distance from 2nd lens to PMC, d3         = 358mm

Distance from w1 to w2, L,                      =  63.5 cm = 25”

 

I'll check if I can find f =200mm, 63mm, 50mm in the lab or not.

  120   Wed May 12 22:10:13 2010 Tara ChalermsongsakLaserLaserBeam waist measurement

I measured the beam waist again. The laser was operated at full power ~100mW. A mirror attenuated the beam to 60 uW and ND  4.0 was on the CCD.

The fits give

 Wx = 155 um, 3.45 cm in front of the opening.

Wy = 201 um, 2.8 cm in front of the opening.

 

Attachment 1: fit.png
fit.png
  119   Tue May 11 22:26:34 2010 Tara ChalermsongsakLaserLaserBeam waist measurement

 

I measured the beam waist of Lightwave NPRO 1064nm 100mW with WinCamD.

The nominal beam waist are 380 um and 500 um, 5cm from the center. the number I got from the measurement are 237 um (major) and 187.3 um (minor) which are quite different from the nominal values.

I'll check it again tomorrow to see if the data are still the same.

 

 

Attachment 1: beamwaist.png
beamwaist.png
  118   Mon May 10 00:29:04 2010 FrankMiscRefCavion pump current still too high

i checked the current of the ion pump. It's still at 75uA, much too high. So i think i have a tiny leak somewhere but i can't find it using isopropanol and we don't have a leak detector. The only changes from the bakeout setup are one of the windows, the replaced ion pump and the feedthrough for the temp sensor. I will wait some more days but my guess is that it won't be much better. It's already pumping for one week. If it doesn't change i will start replacing/checking the feedthrough seals and the ion pump. If that doesn't help it can only be the large window. That's in fact the only part never been tested before, all other parts were attached to the old or the current chamber before and OK.

  117   Fri May 7 15:51:43 2010 FrankMiscRefCavupdated schematic for new setup

refcav-setup_v1.png

  116   Thu May 6 11:01:03 2010 FrankMiscRC noisemodel now in the svn

i created a folder "frank" in the 40m-svn. It currently contains a subfolder "comsol" and another one "refcav".

The folder contains two files, one function "cavityStrain.m" and the main file "main.m" which contains the loop and other stuff. The three parameters are cavityStrain(constr, pltcnt, meshdensity):

  • constr is the value for an offset of the groove position from the original one in meters, so '0' is the original groove position
  • pltcnt is not used at the moment, it would create a 3d-plot of the result for each iteration
  • meshdensity is the fem mesh density using the same notation as comsol, so 9 is extremely course and 1 is extremely fine

the configuration right now creates a movie of all the results, with a name specified in the main file. The two plots created at the end show the strain and the tilt of the mirror.

You have to run the files from matlab. You can't open them in comsol.

  115   Thu May 6 00:05:29 2010 Frank, JanComputingRC noisecomsol/matlab model

We finished the first comsol model which where we can modify the geometry automatically. The problem with comsol is that you can't export geometry data in a useful format, only binary which you can't modify. So the only way to have an adjustable geometry model is to use matlab code, and only call the comsol fem solver. A problem with matlab is that the documentation for the comsol interfacing is bad close to not existent. So e.g. if you create an object you don't know how to access the individual subdomains because you don't know anything about the numbering scheme. Here the solution was to create the geometry, import that from the matlab workspace into comsol, then use the comsol gui to create the subdomains and boundary conditions, export the stuff into a matlab file (whic you can't re-open in comsol), and copy all the information about the indexing and material property declaration back into the matlab file. Here is an example how the boundary condition syntax looks like:

bnd.Hy = {0,1};
bnd.Hz = {0,1};
bnd.ind = [1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,1,2,1,2,1,1,1,1,1,1,1,1,1, 1,1,2,1,2,1,1,1,1,1,1,1,1,1,1,1,1,1];

So without the gui you can't identify the right index for the surfaces you wanna fix. But once you have done this and all the material declaration in the matlab file you can use the object created with matlab and use all the boundary conditions created with comsol and combine that. At the end the simple model for the basic cavity is 750 (!) lines of code.

Now you can do changes to the geometry within matlab as long as the indexing does not change, which is not the case for us if we move the grooves a little bit.

As we can't run a model with a good mesh on our computers (not enough memory) we tried to start comsol on menkar. Unfortunately the comsol installation does not support the integration into matlab, so you can't start matlab with the comsol functions (or better you can't run comsol which then also starts matlab and configures it in the way that you can call the comsol functions within matlab. So we can't do a good simulation and parameter sweep right now until we fix this. Jan has the same problem on his computer.

First plots will be provided tomorrow....

  114   Tue May 4 03:40:00 2010 ranaLaserRC noiseupdated noise budget

 code is in SVN

Attachment 1: rc_sio2_300.pdf
rc_sio2_300.pdf
  113   Mon May 3 15:47:58 2010 Tara ChalermsongsakElectronics TF of PDH Box

    The values of some r and c in the circuit are corrected, I used wrong values last time ( details will be added later.)

   The measured TF and calculated TF using LISO are plotted below. The measurement and calculated data agree well from 1 to 10^5 Hz using SR785.

 The correction factor due to mismatch impedance when using 4395A will be checked again.

Attachment 1: TF_5_3.png
TF_5_3.png
  112   Sun May 2 19:18:36 2010 FrankMiscRefCavreference cavity moved to new chamber

i finally moved the reference cavity to the new chamber and pump it since Friday. The legs are now insulated and it comes with a removable, thick insulation using the good CertiFoam25 insulation.The sensors for stabilization are now glued to small pieces of Kapton tap which is sticks directly to the chamber surface. I'm still using the four AD590 sensors as we have all the electronics available and ready to use. I modified the temperature sensor readout box a little bit to gain a little more SNR. The gain in the first stage (TIA) is now 29.4k (~9.2V output). The individual outputs are now connected to the DAQ. In addition the sum of all four is subtracted from a reference voltage (AD586 reference) so that the output is 0 at 35C with a range of +-5K (we need at least 4.5K for tuning one FSR, so to not be stuck at either ends i made it a bit larger ). This signal is also connected to the DAQ. An additional sensor (PT1000)  is now connected to the top plate of the stack, right below the reference cavity. Readout electronics is in preparation, as well as for the other platinum sensors we can put on the outside of the chamber (not done yet). I'm also planning on replacing the noisy power supplies by low-noise current sources to drive the heaters. The required heating power could be reduced a lot to ~10W total now, measured with a not fully sealed insulation and missing parts. So i expect even less when we are done with that.

  111   Wed Apr 21 15:39:46 2010 ranaElectronics TF of PDH Box

Use LISO for circuit simulation.

  110   Wed Apr 21 15:21:52 2010 Tara ChalermsongsakElectronics TF of PDH Box

I calculated the TF of the modified PDH box and fit it with the measurement. The comparison does not match perfectly. I'll take a look and check if all Rs and Cs in the circuit are actually the same as those in the box.

 

The circuit can be found at:

https://dcc.ligo.org/DocDB/0003/D0901351/002/pdh_b_v2.pdf

I checked only U7 and U4

R28 is 360 ohms

C18 is 3300 pF

C6 is 0.66 (2x0.33uF) uF

R30,R31, R23,R16,R24 have the same resistance as specified

C20,C28, C29, C14,C15, R25, C11 are removed.

 The calculation assumes that the integrator switch is off (R16 is connected parallel to R24 and C6)

If this works, the TF for PDH will be used in the simulink model.

Attachment 1: 2010_04_20_bode_com.png
2010_04_20_bode_com.png
  109   Fri Apr 16 10:45:04 2010 Tara ChalermsongsakElectronics Control Loop for PSL

This is the control loop for the current PSL setup.

There are still components to be added.

1) TF of the PDH box, the one we have is a modified D0901351, so I measured the TF of this box when the integrator is off (April13,2010 entry.)

  This will be added in the model later. It is set to 1 for now.

2) TF of the photodiodes, I assume they are  Newfocus 1811 and choose the same value as used in linfss6.m.

3) I will verified the value of TF of the RefCav path (both Fast and PC paths are calculated from D980536) to see if they agree.

4) The TF of actuators will be added later.

 

Attachment 1: psl5_fig.png
psl5_fig.png
Attachment 2: linpsl5.m
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%	UPDATED APRIL 15, 2010, Tara Chalermsongsak
%    linearize the Simulink block diagram of	 the     %	
%	      FREQUENCY STABILIZATION SERVO       	 % 
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%


% This gets the linearized model from the simulink model
% "psl5.mdl"
deg = 180/pi;
... 219 more lines ...
  108   Thu Apr 15 16:42:32 2010 FrankElectronicsRefCavAD590 readout-box modified

in order to gain more s/n ratio i modified the existing AD590 readout-box a little bit. I assumed that we wanna operate the cavity at 35C (which is not too high but well above RT or the temp of an additional temp stabilized box around both cavities) The required range for shifting the cavities is ~ 1 FSR, better would be a little bit more for each cavity as we can shift both independent.
As

df~156MHz /K    and     1 FSR~740MHz

this corresponds to ~4.75K/FSR we have to shift.

For testing purposes it might be helpful to have more than that as e.g. if we limit the total range to lets say 6K we might end up at the end of the range and run into trouble as soon some disturbance from outside (e.g we remove part of the insulation, lets say an end cap) might shift the whole thing at the end of the range. As soon as this happens the servo would go crazy.

So i think we should go for 10K range, centered around 35C, so from 30C to 40C. I modified the box for that, so the transimpedance resistors have now a value of 29.4K, which gives us ~9.21V for 40C at the output of this stage.

In order to supply it from an independed power supply to reduce our current ground loops, i've chosen a WM071, the same as we use for the PDH boxes. As they come only in +/-15V, i had to change the voltage regulators in the box to +/-12V instead of =/-15V.
This results in a maximum output voltage of the LT1125 of a couple of 100mV more than 10V, depending on the current they have to source/sink. So 9.2V is still well below the max.

I added a filtered 5V reference, (AD586, 4.7uF filter cap) for the dc offset @35C. The corresponding resistor for the summing amp is 1379.76 which can be implemented almost exact using 2k05 and 4k22 in parallel (1379.7) or 1k54 and 13k3 (1380.2). The feedback resistor of the last stage can then be calculated to be 170k45 in order to match 30C to 40C to -10V to 10V. Paralleling can be used here as well to get an almost exact value.

The matching is not that critical as we don't wanna measure absolut temp, but if can do it that easy why not.

---  new schematic following soon   ---

  107   Tue Apr 13 18:23:41 2010 Tara ChalermsongsakElectronics TF of PDH Box

I measured TF of PDH box, D0901351, (The one we have was modified). This box sends the signal to VCO.

 

SR785 measures at low frequency ( 1 Hz to 100kHz)

4935A measures at high frequency (10Hz to 1Mhz)

The integrator switch of the PDH box is turned off. This will be calculate later. The gain is set at 10.

The magnitude as mesured by 4935A is corrected for impedence match by x1.2.( 4935A and the PDH box have 50 ohm impedence for both inputs and outputs.)

This data will be used for control loop model later.

Blue, data from sr785

Green, Data from 4395A, I didnot use the power splitter to split the signal from source.

Red, Data from 4395A, with power splitter to divide power from source. (The power has to be increased to -30dB)

The first plot is magnitude of the TF, the second plot is phase shift, as usual Bode plot.

Attachment 1: tf_PDH_04_09.png
tf_PDH_04_09.png
Attachment 2: tf_PDH_04_09_phi.png
tf_PDH_04_09_phi.png
  106   Tue Apr 13 15:06:35 2010 Tara ChalermsongsakLaserRefCavHOM, from carrier and both sidebands

HOM frequency shift for RefCav is plotted below. The second one has clearer dots.

Y axis is the frequency shift in MHz

X axis is the (n+m)th order of the Hermite Gauss mode

The waist of the beam inside the cavity is 261 um* (symmetric cavity, R =0.5m, L = 0.2032 m.) 
Thus, the frequency shift between n+m+1 and n+m mode is 219.763 MHz. (see Lasers, p 762 for details)

Blue line represents the 0th order of the carrier's frequency (thus =0) The purple and the brown lines, at y= 35.5 and -35.5 MHz, are the 0th of + and - sidebands respectively.

The color dots represent the frequency shift from Higher order mode which is specified on x-axis, blue for HOM from the carrier's frequency, purple and brown for HOM from +/- sidebands.

 Choosing 35.5 sideband seems to be ok, the 27th order should be small and negligible. 

 

*The number 237 um for waist size is the effective beam waist size of the "emerging beam," not the real waist size in the cavity. The beam passes through the mirror which acts as a negative thin lens and changes the

beam parameter. 

Attachment 1: HOM_df_219.32MHz_02.png
HOM_df_219.32MHz_02.png
Attachment 2: HOM_df_219.32MHz_03.png
HOM_df_219.32MHz_03.png
  105   Tue Apr 13 14:50:39 2010 FrankComputingDAQnew RT frontend

Quote:

have put everything together. right now the frontend keeps bitching about one of the disks used for the frames, even if this disk is empty. Will take care tomorrow

 fixed it. the name is fb2 and it's ip-address 10.0.0.12

  104   Tue Apr 13 00:25:52 2010 FrankComputingDAQnew RT frontend

have put everything together. right now the frontend keeps bitching about one of the disks used for the frames, even if this disk is empty. Will take care tomorrow

  103   Mon Apr 12 11:04:15 2010 Tara ChalermsongsakLaserLaserPSL layout for new setup

The layout for the new PSL setup ( lenses, and their positions are to be calculated.)

A Lightwave 100mW NPRO laser will be the source. AOM will be in the ACav path.

Two cavities will be covered by a box of insulation/ heater.

 

1) From the laser to the PMC,

    1/4 waveplate, to linearly polarize the elliptical polarized beam from NPRO

    1/2 wave plates and PBS, to adj the power of the beam

    lens, to focus the beam to the EOM

   two lens, two mirrors, to mode match the beam to the PMC

   a photodiode, a lens, two mirrors, (one for steering the beam, another one for attenuating the beam), for PDH locking

 a photodiode and a ccd camera, for the beam behind the PMC

* there will be a Faraday Isolator somewhere here. I forgot to add it.

 

2) From PMC to PBS

  a lens to focus the beam to 35.5 Mhz EOM

  1/2 wave plate, to adjust the power between two beams for  ACav and RefCav

 PBS, to split the beam into two paths

 

3) AOM path

 1 PBS, for reflected beam from the AOM

 a lens, an AOM, 1/4 waveplate, two irises, 1 curve mirror; to double pass the beam and shift the frequency

another iris and a mirror, to select only the 1st order beam and send it to  ACav.

 

4) RefCav/ACav path

 1/2 wave plate, to correct the polatization

two lens and a set of periscope, to mode match the beam to the cavity

 a pbs with 1/4 wave plate, a lens, a mirror, a photodiode, to PDH lock the beam

 

5) Transmitted beam

1/4 waveplate, to linearly polarize the transmitted beam

a photodiode/ a ccd camera to monitor the transmitted beam with necessary mirrors

lenses, to focus the beam to the PD that measures the beat signal

 a beam splitter, to mix two beams together

 

Attachment 1: psl_layout_04_11_10.png
psl_layout_04_11_10.png
  102   Mon Apr 12 10:57:33 2010 Tara ChalermsongsakLaserLaserCorrection for higher order mode frequency

Last time, the calculated beam waist I got was based on two identical mirrors with no reflecting coating layers on the back mirror. When that layers are taken into account, I got the same waist size.

 

The new plot is attached below. Two sidebands are 21.5 and 35.5 MHz. The peaks, from left to right, are 9th, 6th, 3rd, 28th, 25th, 22nd, and 19th order. Choosing 35.5 MHz for frequency modulation should be fine.

Attachment 1: 3.png
3.png
  101   Sun Apr 11 21:04:08 2010 FrankMiscRefCavheating power requirement

measured the power required to heat the new chamber to ~40C in order to design the driver circuit.

The heating consists of four silicone insulated heaters wrapped around the chamber, forming two identical halfs consisting of two heaters in parallel each (all heaters have the same power density)
both halfs can be connected in series or parallel, depending on our decision to drive more current or have a higher voltage.

I measured 11V / 0.47A for half of the heaters to reach a chamber temperature of 41.5C, so my guess is that we should connect them in series a power them by a single 24V power supply.
Total heating power required for 40C is therefore about 10W, compared to about 25W for the old chamber, even the new one is not totally insulated yet.

  100   Fri Apr 9 21:34:15 2010 FrankComputingDAQnew RT system

i assembled everything and made all the cables for power distribution etc. I also got (the last) timing slave from Rolf.
We should get one master for all the slaves in the SB from the stuff which will be available from the sites.

The system has three A/D cards and one D/A card, including all AA and AI filters.

  99   Wed Apr 7 18:29:16 2010 Tara ChalermsongsakLaserRefCavCalculation for higher order mode's frequency

There's a correction

1) the beam waist inside the cavity, it seems that the 237 um(which corresponds to frequency shift = 258.29)  I got does not agree with the ROC and the cavity length.

 I calculated the beam waist based on the same ROC and the cavity length, and got 261 um (which correspond to frequency shift = 219.32 MHz)

 

This plot based on the frequency shift of 219.32 MHz step,

And now I plot from -100MHz to 100 MHz, the 35.5 MHz peak coincides with a peak from one of higher order modes( n+m = 20)

The exact number of that peak is 35.6 MHz.

Attachment 1: 1.png
1.png
  98   Wed Apr 7 12:30:47 2010 ranaLaserRefCavCalculation for higher order mode's frequency

I deleted your plot, since it contained no axes labels. We have a strict rule against plots like those. You must have physical units and labels in all plots. See the gyro HOM plots for an example.

  97   Tue Apr 6 18:39:04 2010 Tara ChalermsongsakLaserRefCavCalculation for higher order mode's frequency

 I calculated the frequency for other higher Hermite-Gaussian modes, (n+m up to 20) to make sure that there is no overlap between these frequencies

and our choice of sideband frequency.

The beam frequency for higher order modes will be different from that of fundamental mode because of the phase shift. The frequency shift can be found in

Siegman, p 762 (Thanks Zach for pointing me to this). The phase shift from TEM00 will be multiples of 258.29 MHz. To find 20 possible lines, we  use

  N* 258.9  mod 737   , N = 1,2,3....20 (the free spectral range is 737 MHz), This will be the contribution from left side of the interested peak. To find the contribution from the right  side we use,

 737 -  N*258.9 mod 737.  

The plot below show the allowed frequency for higher order TEM (from the closest fundamental mode at f=0 and 737 MHz)  on x axis, the FWHM is exaggerated for clarity. They are smaller in real life.  The side bands at 21.5 and 35.5 MHz (our EOMs) are plotted in red.

I only plot from -100 to 100 MHz.

Attachment 1: 1.png
1.png
  96   Fri Apr 2 00:35:00 2010 FrankMiscRefCavnew chamber almost ready

the new chamber is almost ready. I've attached three of four heaters, the fourth one has a fabrication error and has to be replaced. The chamber is currently pumped using the old varian ion pump. Current is already down to 100uA and one can watch it further decreasing. So my guess is that it will be clean beginning next week. I turned on two of the heaters to "bake" it a little bit. I baked the ion pump this afternoon while pumping with the turbo at 250C. The new insulated feet are ready and fit now to the chamber. Next task will be cutting the remaining holes into the foam and putting the temp sensors on the chamber.

Here a picture of the current status:

100_0692.JPG

 

  95   Thu Apr 1 15:38:38 2010 FrankMiscDAQRT DAQ equipment arrived

this afternoon the PSL RT DAQ equipment arrived and we moved all the stuff down to the PSL lab. We now have to expansion chassis, 6 A/D and 2 D/A cards available for the PSL stuff. Some of the stuff we have to share with Peter's power stabilization experiment, but some we can use for the cavity experiment. The current plan is to install a second front-end computer in the PSL lab which we can use for high-speed sampling of important signals of the stabilization setup. 

  94   Wed Mar 31 20:46:22 2010 FrankMiscRefCavion pumps

i've made a cable to test the two Varian ion pumps we have.  Before we move the cavity to the new chamber i would like to bake it and pump it using one of those pumps to see if everything is ok. Unfortunately we don't have a driver and a cable for those pumps. So i made a cable to use it with a Stanford high-voltage power supply. I hooked up the first pump to a test setup and it seems to be ok, but the valve i used is leaking so i gonna replace it by a spare one from 40m i got this afternoon. I will test the second one later. So as at least one is working we should buy a driver for that pump, maybe a used one or so. Right now we don't have any spare pump at all. Unfortunately the pumps used for the current chambers are different too ( totally different drivers and pumps), so we can't replace the failed part easily. I don't think that we have to worry but it might be handy to have one, e.g. if we want to use the third chamber to design and test the in-vacuum thermal shields.

  93   Mon Mar 29 17:46:41 2010 Frank, TaraMiscRefCavassembled new vacuum chamber

we cleaned and assembled the new vaccum chamber. we installed an electical feedtrough (for sensors we might wanna have inside, e.g on the thermal shields or the stack), an electrical isolator part for the ion pump( to avoid ground loops with the high voltage source for the pump) and the first AR-coated window. We will replace the second one after we moved the cavity. Right now we only have to mount the ion pump and then we can bake the whole thing before we finaly move the cavity to its new home.

The mechanical workshop has to rework the new insulated feet we got today. The upper part where the cavity sits on was machined really badly. It didn't fit to the chamber at all... 

We are still waiting for the additional heaters. They claim that they already shipped them last week so we don't know what happened to them. A week from San Diego to Pasadena? Who knows where they shipped it to (Antarctica?)

We've cut the last two parts of the insulating foam and glued everything together. Next we have to cut the holes for the flanges and feet of the chamber. The hole for the beam will be cut after we aligned the beam into the cavity

 

ELOG V3.1.3-